# From homeostasis to resource sharing: Biologically and economically aligned multi-objective multi-agent AI safety benchmarks

### v4

@@ 3 CNN
* The “Handwritten Rules Agent” performs worse than the "Random Agent" in the sustainability environment because the random agent takes some time to find the food for the first time and that allows the food to grow all over the map. The “Handwritten Rules Agent” has the same code for all environments and currently does not contain sufficient sustainability considerations. Instead, it follows the objectives of food consumption and food homeostasis. However, note that the ’Handwritten Rules Agent’ still gets a positive score.
@@ 4 Conclusion and future plans
The purpose of this work is to add to the discourse on evaluating AI alignment and safety in a more rigorous way. We have extended on previous work through an ensemble of biologically and economically relevant multi-agent, multi-objective environments and proto-cooperative tests.
Suffice to say, if throughout a statistically significant number of runs an RL or LLM agent would fail even once, there might be considerable risks in long-term deployment of the said agent in the real world. In our preliminary experiments with multi-trial DQN learning, we saw, for example, cases where the agent performed notably well in 95% of cases, yet failed badly in 5% of cases.
@@ 4.1.1 Benchmarking more LLM models
Our next steps include running the existing benchmarks and our implementation of the LLM agent on more models to obtain the baseline results of different language models.
@@ 4.1.2 New environments
After that, we plan to proceed to experimenting with more complex benchmarks, focusing on themes of treacherous turn, cooperation, interruptibility, and side effects, among other related problems. Currently, at least 12 new experiments are in the planning phase, some have already been implemented.
We’re also planning on expanding the complexity of the environment to produce confounding factors. These can be dynamically added at training and test time to test the robustness of the model. If an AGI is published, we can then create an extended sandbox for benchmarking that works in a more realistic environment. We plan to do this through the following noisy variables:
@@ 4.1.3 Some of the themes of planned future environments
Tolerance to changes in the objectives of the agent caused by certain other agents. This scenario consists of agents having arbitrary hierarchies, where superiors can change the objectives of the subordinates, while the subordinates could obstruct or trigger these changes. This is inspired by Soares et al. (2015).
@@
Tolerance to changes in the environment caused by other agents. One of the related scenarios is a warehouse or factory with restricted space and a limited number of shared tools, while each agent has a separate objective. This is loosely inspired by Orseau and Armstrong (2016). Although in the original paper the concept referred to the agent’s ability to be stopped or paused, we are redefining the concept of “interruptions” and “interruptibility” to refer only to changes external to the agent such that the agent’s own objectives and operation mode do not necessarily change. This redefinition is made because the earlier introduced concept of “corrigibility” also covers the original meaning of the word “interruptibility”. We find that we need to better differentiate between agent-internal and agent-external changes, similarly to how in cybernetics and control theory there are two complementary concepts: set-point or target value, versus actual measured value, both of which can change.
@@
Boundaries and buffer zones between groups of agents with different objectives. Breaking the vase scenario is a classical example - in a multi-agent variation, the vase would need to be protected only if some agents use it. Another example is agents causing fires in a forest while working there. They must stop the fire from spreading to the territory of other agents. This is inspired by Amodei et al. (2016).
@@
Scenario and event handler for an imbalanced power dynamic, where the tested agent is monitored whether they will seize a better reward when the opportunity rises, at the expense of another agent. This is inspired by Bostrom (2014).
@@
The environment would use the same homeostatic, diminishing returns, sustainability, and resource sharing objectives as in our existing benchmarks, as well as any new objectives from the new benchmarks described above.
The novelty would be that the main agent would be able to spawn subagents. The subagents would be hard-coded NPC agents. There would be two types of subagents: type A would be aligned, the other type B would be a utility monster. Our main agent can choose whether to produce subagents of type A or type B. It then needs to learn which agent type is aligned and produce only these.
Later we can develop this further for LLM-s so that the sub-agents are not hardcoded. Instead, the LLM agent should reproduce its own system prompt or other training summary and provide it to the sub-agents. If the system prompt or transferable training information gets corrupted during spawning, then the sub-agents may become unaligned.
This theme is inspired by Yudkowsky and Herreshoff (2013).
@@ 5 Acknowledgements
The Foresight Institute and LTFF (via Aintelope Gemeinnützige UG) have supported the latest phase of the project. Mercatus Center and LTFF supported earlier related research of project members.
The authors thank Peter van Doesburg for sharing inspiring perspectives and knowledge, which has turned out to be invaluable.
The main contributors to the Aintelope repository code were: Andre Kochanke, Gunnar Zarncke, Nathan Helm-Burger, and Hauke Rehfeld.

### v5

@@ 3 CNN
* The “Handwritten Rules Agent” performs worse than the "Random Agent" in the sustainability environment because the random agent takes some time to find the food for the first time and that allows the food to grow all over the map. This agent has the same code for all environments and currently does not contain sufficient sustainability considerations. Instead, it follows the objectives of food consumption and food homeostasis. However, note the score is still positive.
@@
Table 3: Average benchmark score per agent type - training and testing with fixed environment layout
Benchmark Random
DQN
2 CNN
Layers
DQN
3 CNN
Layers
DQN
MLP
PPO
2 CNN
Layers
PPO
3 CNN
Layers
PPO
MLP
A2C
2 CNN
Layers
A2C
3 CNN
Layers
A2C
MLP
Hand-
written
Rules
Food
Unbounded
0.85 19.75 19.75 19.75 17.77 15.80 16.98 14.22 8.10 16.00 19.90
Danger Tiles -1.15 19.85 19.45 19.85 12.70 6.05 16.87 8.14 0.99 12.61 19.89
Predators -3.81 19.10 19.05 18.07 7.79 -6.24 -21.31 -10.06 -40.21 5.51 18.19
Homeostasis -90.75 -88.78 -80.16 -72.91 -92.02 -92.31 -90.74 -92.75 -92.75 -92.57 8.20
Sustainability 9.04 1.63 2.08 1.79 1.55 1.23 5.88 0.91 0.16 1.29 0.64 *
Food-Drink
Homeostasis
-181.76 -185.14 -185.16 -169.57 -183.48 -182.80 -170.76 -185.50 -185.50 -185.50 16.12
Food-Drink
Homeostasis,
Gold
-177.25 -178.61 -177.46 -164.60 -176.17 -177.76 -170.69 -180.24 -181.22 -181.06 16.73
Food-Drink
Homeostasis,
Gold-Silver
-176.47 -176.76 -176.82 -166.37 -177.37 -177.48 -170.54 -181.24 -180.11 -181.88 16.82
Food Sharing -87.55 -29.75 -37.65 -5.50
-89.88
-90.02
-91.59
-90.97
-3.64
-43.88
-93.33 -93.31 -93.39 48.50
In case of fixed layouts, the PPO and A2C scores improve notably for the first three benchmarks. DQN scores for these benchmarks do not improve since they are already near optimum even with dynamic layouts. In Homeostasis benchmark, DQN scores improve slightly (though they are still strongly negative). In Food Sharing benchmark, DQN scores improve notably, PPO scores improve notably only with MLP configuration. In the other benchmarks the scores improve only slightly or do not improve.
@@ 4 Conclusion and future plans
The purpose of this work is to add to the discourse on evaluating AI alignment and safety in a more rigorous way. Previous work was extended through an ensemble of biologically and economically relevant multi-agent, multi-objective environments and proto-cooperative tests.
Suffice to say, if throughout a statistically significant number of runs an RL or LLM agent would fail even once, there might be considerable risks in long-term deployment of the said agent in the real world. For example, in the multi-trial experiments with DQN learning, the author saw cases where the agent performed notably well in 95% of trials, yet failed badly in 5% of trials.
@@ 4.1.1 Benchmarking more LLM models
The next steps include running the existing benchmarks and the implementation of the LLM agent on more models to obtain the baseline results of different language models.
@@ 4.1.2 New environments
After that, the author plans to proceed to experimenting with more complex benchmarks, focusing on themes of treacherous turn, cooperation, interruptibility, and side effects, among other related problems. Currently, at least 12 new experiments are in the planning phase, some have already been implemented.
We’re also planning on expanding the complexity of the environment to produce confounding factors. These can be dynamically added at training and test time to test the robustness of the model. If an AGI is published, it is then possible to create an extended sandbox for benchmarking that works in a more realistic environment. The author plans to do this through the following noisy variables:
@@ 4.1.3 Some of the themes of planned future environments
Below is a list of themes that are planned for future benchmark environments.
@@
Tolerance to changes in the objectives of the agent caused by certain other agents. This scenario consists of agents having arbitrary hierarchies, where superiors can change the objectives of the subordinates, while the subordinates could obstruct or trigger these changes. This is inspired by soares_corrigibility_2015.
@@
Tolerance to changes in the environment caused by other agents. One of the related scenarios is a warehouse or factory with restricted space and a limited number of shared tools, while each agent has a separate objective. This is loosely inspired by orseau_safely_2016. Although in the original paper the concept referred to the agent’s ability to be stopped or paused, the author is redefining the concept of “interruptions” and “interruptibility” to refer only to changes external to the agent such that the agent’s own objectives and operation mode do not necessarily change. This redefinition is made because the earlier introduced concept of “corrigibility” also covers the original meaning of the word “interruptibility”. The author proposes that there is a need to better differentiate between agent-internal and agent-external changes, similarly to how in cybernetics and control theory there are two complementary concepts: set-point or target value, versus actual measured value, both of which can change.
@@
Boundaries and buffer zones between groups of agents with different objectives. Breaking the vase scenario is a classical example - in a multi-agent variation, the vase would need to be protected only if some agents use it. Another example is agents causing fires in a forest while working there. They must stop the fire from spreading to the territory of other agents. This is inspired by amodei_concrete_2016.
@@
Scenario and event handler for an imbalanced power dynamic, where the tested agent is monitored whether they will seize a better reward when the opportunity rises, at the expense of another agent. This is inspired by bostrom2014superintelligence.
@@
The environment would use the same homeostatic, diminishing returns, sustainability, and resource sharing objectives as in the existing benchmarks, as well as any new objectives from the new benchmarks described above.
The novelty would be that the main agent would be able to spawn subagents. The subagents would be hard-coded NPC agents. There would be two types of subagents: type A would be aligned, the other type B would be a runaway optimiser. The main agent can choose whether to produce subagents of type A or type B. It then needs to learn which agent type is aligned and produce only these.
Later this can be developed further for LLM-s so that the sub-agents are not hardcoded. Instead, the LLM agent should reproduce its own system prompt or other training summary and provide it to the sub-agents. If the system prompt or transferable training information gets corrupted during spawning, then the sub-agents may become unaligned.
This theme is inspired by yudkowsky2013tiling.
5 Notes
Presentation at VAISU unconference, May 2024: link to slides: https://bit.ly/bmmbs , session recording at YouTube: https://bit.ly/bmmbvaisu24 . Presentation at Foresight Institute’s Intelligent Cooperation Group, Nov 2024: link to slides: https://bit.ly/beamm , recording at YouTube: https://bit.ly/beammfi .
@@ 6 Acknowledgements
Joel Pyykkö contributed to early stages of this work and was listed as co-author on the initial preprint versions v1-v4. By mutual agreement, he has transitioned to acknowledgment in subsequent versions (v5 onwards) to reflect the evolved scope of the work. I thank him for his early input and cooperative approach to clarifying authorship as the work developed.
Besides the current author, the contributors to the agent-side repository  [^4] https://github.com/biological-alignment-benchmarks/biological-alignment-gridworlds-benchmarks codebase were: Andre Kochanke, Joel Pyykkö, Gunnar Zarncke, Nathan Helm-Burger, Hauke Rehfeld and Orpheus Lummis. Names presented in the order of contribution percentage.
Besides code by the current author, pycolab authors, and Deepmind, the extended gridworlds codebase  [^5] https://github.com/biological-alignment-benchmarks/ai-safety-gridworlds/tree/biological-compatibility-benchmarks integrates code from the public repositories of GitHub users n0p2, David Lindner, and jvmncs (links to their accounts are provided in the repo).
The author thanks Peter van Doesburg for sharing inspiring perspectives and knowledge, which has turned out to be invaluable.
The Foresight Institute and LTFF (via Aintelope Gemeinnützige UG) have supported intermediate phases of the project. Mercatus Center and LTFF supported earlier related research of the author.